OpenStudy (anonymous):

Write the equation of the line that is perpendicular to y= 1/4x and passes through the point (2,-5)

OpenStudy (psymon):

So you'll first be finding the slope and then after you find the slope, you use the point-slope form equation. It also asks for a perpendicular slope. A perpendicular slope is when you take the negative reciprocal of your current slope. So if you had a slope of 2, the perepndicular slope would be -1/2. That being said, the initial equation ius y=1/4x, which is the in the form of y = mx + b. If m is the slope, then in your equation, 1/4 is the slope. So knowing that the perpendicular slope means takign the negative reciprocal, do you know the point-slope form in order to do the rest of the work?

OpenStudy (anonymous):

no i dont

OpenStudy (texaschic101):

is that the total equation....y = 1/4x ??

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

hold on.....@gypsy1274 .....can you take a look at this

OpenStudy (texaschic101):

@gypsy1274

OpenStudy (texaschic101):

@campbell_st ...help

OpenStudy (texaschic101):

do you do it the same way even though there is no y intercept psymon ?

OpenStudy (psymon):

well, point slope form is y - y1 = m(x-x1) So the idea is to know what your perpendicular slope is and then just plug in numbers. And the y-intercept isnt important, just that number in front of x.

OpenStudy (texaschic101):

ok......wasn't sure....thanks

OpenStudy (psymon):

Mhm. So if the slope is 1/4, then the perpendicular slope would be flipping it and making it negative (negative reciprocal). Then that value becomes your new "m" in the point slope form equation I listed. The y1 and x1 are just the cooridnates of the point you were given.

OpenStudy (campbell_st):

An easy way to do this question is slope of the given line is m = 1/4 the perpendicular has a slope that is the negative reciprocal m = -4 the slope intercept form is y = mx + b substitute you slope. the substitute the information from your point (2, -5) which is x = 2 and y = -5 to find the value of b.

OpenStudy (campbell_st):

oops should read slope for the perpendicular... into y = mx + b

OpenStudy (anonymous):

so how would you write out the equation ? -4+-5 ?

OpenStudy (campbell_st):

y = -4x + b substitute x = 2 and y = -5 -5 = -4*(2) + b solve for b

OpenStudy (anonymous):

-1.6 ?

OpenStudy (campbell_st):

nope

OpenStudy (texaschic101):

-5 = -8 + b -5 + 8 = b can you take it from here ?

OpenStudy (psymon):

y+5 = -4(x-2), then you multiply and solve from there. That was just what I plugged into the point slop form equation. You would get what they're getting as well.

OpenStudy (texaschic101):

thats true

OpenStudy (anonymous):

i add -5 and 8 ?

OpenStudy (texaschic101):

that is your y intercept......so yes

OpenStudy (texaschic101):

y = -4x + b (b is your y intercept)

OpenStudy (anonymous):

so 3 is the y

OpenStudy (texaschic101):

no...3 is the b

OpenStudy (texaschic101):

your answer is y = -4x + 3

OpenStudy (texaschic101):

or psymon's way... (2,-5) y - y1 = m(x - x1) y - (-5) = -4(x - 2) y + 5 = -4x + 8 y = -4x + 8 - 5 y = -4x + 3

OpenStudy (anonymous):

its wrong ...

OpenStudy (texaschic101):

huh.....really ?

OpenStudy (anonymous):

yea i just tried it

OpenStudy (texaschic101):

@campbell_st ......she said it was wrong

OpenStudy (campbell_st):

so y = -4x + 3 is wrong, what about 4x + y = 3 or y + 5 = -4(x - 2) any of those are correct.

OpenStudy (texaschic101):

true.....it could be in Ax + By = C form.....or slope intercept form

OpenStudy (anonymous):

there all wrong ?

OpenStudy (anonymous):

@campbell_st

OpenStudy (texaschic101):

what are we doing wrong ?

OpenStudy (campbell_st):

@xoxox123 ok... can you now go and check the question is correctly typed... as based on the information provided... the solutions are correct...

OpenStudy (anonymous):

Yes I'm positive the problem is right!! looked over it @campbell_st

OpenStudy (campbell_st):

ok.... the text answer or website answer is incorrect...

OpenStudy (anonymous):

no this website makes sure you get it right so your problem is wrong

OpenStudy (texaschic101):

I am gonna go with campbell......it looks correct to me

OpenStudy (anonymous):

it looks correct to me too but, i dont know why its counting it wrong

OpenStudy (texaschic101):

is their answer choices ?

OpenStudy (anonymous):

nope. written response

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

yes......I don't know what to say

OpenStudy (texaschic101):

@phi .......please look at this

OpenStudy (anonymous):

hes offline lol

OpenStudy (texaschic101):

oops

OpenStudy (campbell_st):

ok...here is a graph of the 2 lines, with an angle shown. it shows your line from the question, the point and the angle between the lines as 90, which means perpendicular...

OpenStudy (campbell_st):

so what solution did they give you...?

OpenStudy (anonymous):

Solution ?

OpenStudy (campbell_st):

yes... you said our answer was wrong... so what solution did your website give..?

OpenStudy (anonymous):

It doesn't give you an answer when you get it wrong.. it just tells you its wrong then you have to try it again

OpenStudy (campbell_st):

lol... unless the site is wrong.... then you'll enter wrong answer till infinity and beyond.......good luck...

OpenStudy (anonymous):

the website isn't wrong? How is it wrong if I've been working on it for a month and have been entering answers in it? that i get right! so you made a mistake in your answer

OpenStudy (texaschic101):

I just don't see the mistake......I would post this question again and see what other people say

OpenStudy (texaschic101):

good luck :)

OpenStudy (texaschic101):

I think they will say the same thing

OpenStudy (anonymous):

we shall see ..

OpenStudy (campbell_st):

ok.... so you have an algebraic solution to your problem, yo have a graphical solution to your problem, 3 unrelated people get the same solution using the information you have posted... and in every case the solution to the information provided is y = -4x + 3 so perhaps posting the question again, going and reviewing your work on perpendicular lines and equations of straight lines may help you get a solution. Or perhaps the website as mis-typed the answer.